OpenStudy (mayankdevnani):

A ball is dropped on to the floor from a height of 10 m. It reboundes to a height of 2.5 m. If the ball is in contact with the floor for 0.01 seconds, what is the average acceleration during contact? a) \[1400 m/s^2\] b) \[2100 m/s^2\] c) \[700 m/s^2\] d) \[2800 m/s^2\]

OpenStudy (mayankdevnani):

@amistre64 @hba @asnaseer @AccessDenied @jhonyy9 @jazy @ganeshie8

OpenStudy (anonymous):

The acceleration of a body is defined as the change of its velocity during an interval of time divided by the interval of time. Expressed algebraically, a = (Vo - Vf)/t where a = the acceleration, Vo = the initial velocity, Vf = the final velocity and t = the interval of time.

OpenStudy (anonymous):

From the basic a = (Vo - Vf) we derive the first as ......................................t ..............................Vf = Vo + at (the acceleration is assumed constant)

OpenStudy (anonymous):

The second defines the relationship between the distance traveled by a body exposed to a constant acceleration. The distance traveled over an interval of time derives from the product of the average velocity during the time interval and the time interval. The average velocity is (Vo + Vf)/2 and the time interval "t" which yields d = (Vo + Vf)t/2. Substituting the Vf derived above, d = [Vo + (Vo + at)]t/2 resulting in ...............................d = Vo(t) + a(t^2) (quite often, s is used in place of d) ......................................................2

OpenStudy (anonymous):

The third combines the first two by eliminating the time interval "t". Therefore, a(d) = (Vo - Vf)(Vo + Vf)t = (1/2)(Vf - Vo)(Vf + Vo) resulting in .....................................t.............2 ..............................Vf^2 = Vo^2 + 2ad

OpenStudy (mayankdevnani):

thnx.... @tictac4456

OpenStudy (anonymous):

no problem do u go to connexus

OpenStudy (mayankdevnani):

is it right?? While falling down, initial velocity = 0, final velocity = v, acceleration due to gravity = 9.8 m/s 2 , height = 10 m The velocity of the ball with which it hits the ground can be found using, v 2 = u 2 + 2as => v 2 = 0 + 2×9.8×10 => v = 14 m/s (downward) While it moves up, initial velocity = u, final velocity = 0, acceleration due to gravity = -9.8 m/s2 [this time the velocity and acceleration are oppositely directed so we have the negative sign] , height = 2.5 m The velocity of the ball with which it leaves the ground can be found using, v 2 = u 2 + 2as => 0 = u 2 - 2×9.8×2.5 => u = 7 m/s (upward) So, Change in velocity = 7 m/s (upward) – 14 m/s (downward) => Δv = 7 m/s + 14 m/s [both in upward direction] => Δv = 21 m/s So, acceleration is, a = Δv/Δt = 21/0.01 = 2100 m/s 2

OpenStudy (anonymous):

yup

OpenStudy (mayankdevnani):

@tictac4456

OpenStudy (anonymous):

thats correct

OpenStudy (mayankdevnani):

thnx..

OpenStudy (mayankdevnani):

@satellite73 help me:)

OpenStudy (mayankdevnani):

@amistre64

OpenStudy (anonymous):

hey ur work is correct but the answer is wrong

OpenStudy (anonymous):

the answer is A

OpenStudy (mayankdevnani):

howz that??

OpenStudy (anonymous):

oh im srry never mind i didnt see the seconds ur answer is correct

OpenStudy (anonymous):

im 100% sure its right

OpenStudy (mayankdevnani):

ohkkk!!

OpenStudy (asnaseer):

I agree - your work and answer is correct

OpenStudy (mayankdevnani):

thnx...

OpenStudy (asnaseer):

yw :)

OpenStudy (anonymous):

do ya take entrepuener class

OpenStudy (mayankdevnani):

what do you mean??

OpenStudy (anonymous):

other than physics do u take entrepuener class

OpenStudy (anonymous):

its like a business class

OpenStudy (mayankdevnani):

no!! never!!! why??

OpenStudy (anonymous):

haha i kinda needed help

OpenStudy (anonymous):

but never mind

OpenStudy (mayankdevnani):

plz you can!!!

OpenStudy (anonymous):

ok lol but if u dont knw it u dont have to help k

OpenStudy (mayankdevnani):

k

OpenStudy (anonymous):

what is gross profit? A)the money left over after the cost of making a product or providing a service B) the cost of labor to produce a product C) the difference between a product's price and the revenue of the company D) a negative net profit, in which company experiences a loss

OpenStudy (mayankdevnani):

(finance) the net sales minus the cost of goods and services sold.

OpenStudy (anonymous):

thnkz lol

OpenStudy (anonymous):

i gave u a medal right

OpenStudy (mayankdevnani):

right!!!

OpenStudy (anonymous):

cool so did u graduate? or u still in high skool

OpenStudy (mayankdevnani):

high school!!!

OpenStudy (anonymous):

what grade

OpenStudy (mayankdevnani):

9 class

OpenStudy (mayankdevnani):

ok @tictac4456

OpenStudy (anonymous):

me2 lol

OpenStudy (anonymous):

what skool?